OpenStudy (anonymous):

Find all real solutions of the following equation log 3 x +log 3 (x-24) =4

OpenStudy (anonymous):

\[\log_{3}x+\log_{3}(x-24)=4 \]

OpenStudy (anonymous):

my answer is no real solutions

OpenStudy (anonymous):

\[\log_3(x^2-24x)=4\] \[x^2-24x=3^4=81\] \[(x+12)^2+81+144=225\] \[x+12=\pm\sqrt{225}\] \[x+12=15\] \[x=3\]

OpenStudy (anonymous):

i came up iwth -27,3 and so neither of these would suit the equation right?

OpenStudy (anonymous):

yes that is right. you cannot use 3 or -27

OpenStudy (anonymous):

k thanx